OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)? (x - 2)2 + (y + 3)2 = 5 (x + 2)2 + (y - 3)2 = 25 (x - 2)2 + (y + 3)2 = 25 (x - 2)2 - (y + 3)2 = 5

OpenStudy (anonymous):

Ok, so @Joscelyna, remember when writing your solutions to include a (^) to show that you are squaring something. Your first answer, you wrote as (x - 2)2 + (y + 3)2 = 5, but you mean to write (x - 2)^2 + (y + 3)^2 = 5 This is important for there is a big difference between the two and you can confuse those helping you in the future. That being said, do you know the general form of an equation for a circle?

OpenStudy (anonymous):

I don't even remember, gimme one sec here.

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2=r^2

OpenStudy (anonymous):

So I'm going to draw you where the center of the circle is ok? 

OpenStudy (anonymous):

And it says that is passes through the point (-2, 0). Like so,

OpenStudy (anonymous):